OpenStudy (anonymous):

what is the best equation to solve The perimeter of a rectangle is 96 ft. The ratio of the length to the width is 7: 5. What are the dimensions of the rectangle

OpenStudy (nirmalnema):

let the dimension be x so the new ratio is 7x:5x 2(7x/12 + 5x/12)=96 24x/12=96 x=96*12 / 24 x=48 length = 28 breadth = 20